# Rare-to-Frequent: Unlocking Compositional Generation Power of Diffusion Models on Rare Concepts with LLM Guidance

### v2

@@ Abstract
State-of-the-art text-to-image (T2I) diffusion models often struggle to generate rare compositions of concepts, e.g., objects with unusual attributes. In this paper, we show that the compositional generation power of diffusion models on such rare concepts can be significantly enhanced by the Large Language Model (LLM) guidance. We start with empirical and theoretical analysis, demonstrating that exposing frequent concepts relevant to the target rare concepts during the diffusion sampling process yields more accurate concept composition. Based on this, we propose a training-free approach, R2F, that plans and executes the overall rare-to-frequent concept guidance throughout the diffusion inference by leveraging the abundant semantic knowledge in LLMs. Our framework is flexible across any pre-trained diffusion models and LLMs, and can be seamlessly integrated with the region-guided diffusion approaches. Extensive experiments on three datasets, including our newly proposed benchmark, RareBench, containing various prompts with rare compositions of concepts, R2F significantly surpasses existing models including SD3.0 and FLUX by up to $28.1\%p$ in T2I alignment. Code is available at https://github.com/krafton-ai/Rare-to-Frequent.
Figure 1: Generated images from prompts with rare compositions of concepts ($=$attribute$+$object; highlighted in red). These objects possess attributes not typically associated with them, making such combinations difficult to observe. While state-of-the-art pre-trained and LLM-grounded text-to-image diffusion models, SD3.0 (Esser et al., 2024), FLUX (BlackForestLabs, 2024), and RPG (Yang et al., 2024), struggle to generate such concepts, our training-free approach, R2F, exhibits superior results.
@@ 1 Introduction
Recent advancements in text-to-image (T2I) diffusion models have achieved unprecedented success in generating highly realistic and diverse images (Zhang et al., 2023a, Saharia et al., 2022). However, these models often struggle to accurately generate images from rare and complex prompts (Samuel et al., 2024). Accordingly, numerous studies have focused on enhancing their T2I alignment performance (Rassin et al., 2024). Recently, several approaches have further tried to ground large language models (LLMs) into diffusion models, so-called LLM-grounded diffusion models, showing state-of-the-art results by effectively leveraging LLMs’ knowledge in diffusion inference.
@@ 2.3 LLM-grounded Diffusion
LLMs have shown promising abilities in language comprehension (Achiam et al., 2023, Zhao et al., 2023, Wang et al., 2024a). Powered by this advance, recent works have attempted to ground LLMs into diffusion models to provide prior knowledge and conditions for compositional image generation. LayoutGPT (Feng et al., 2024), LMD (Lian et al., 2023), and CompAgent (Wang et al., 2024c) use LLMs to decompose a given prompt into sub-prompts per object and extract their corresponding bounding boxes. RPG (Yang et al., 2024) further adopts recaptioning and planning for complementary regional diffusion. ELLA (Hu et al., 2024) use LLMs to dynamically extract timestep-dependent conditions from intricate prompts. Some work utilizes LLMs in iterative image editing (Wu et al., 2024, Gani et al., 2023). While these LLM-grounded diffusion approaches have succeeded in spatial compositions, they still face challenges in non-spatial compositions for rare concepts.
@@ 3.1 Compositional Text-to-image Generation for Rare Concepts
Theoretical Motivation. To provide an insight into how relevant frequent concepts help the rare concept composition, we consider a simple setting where two texts are given: one involving a rare concept ${\bm{c}}_{R}$ (e.g.,“furry frog”), and another involving a frequent one ${\bm{c}}_{F}$ (e.g., “furry dog”). Let ${\bm{x}}\in\mathbb{R}^{2}$ be an image representation, where the first and second dimension represent the attribute (e.g., “furry”) and object (e.g., “animal”), respectively. We assume the ground truth conditional distributions $p_{\text{data}}({\bm{x}}|{\bm{c}}_{R})$ and $p_{\text{data}}({\bm{x}}|{\bm{c}}_{F})$ follow Gaussian distributions ${\mathcal{N}}({\bm{\mu}}_{R},{\bm{\Sigma}}_{R})$ and ${\mathcal{N}}({\bm{\mu}}_{F},{\bm{\Sigma}}_{F})$, with ${\bm{\Sigma}}_{R}={\bm{\Sigma}}_{F}={\bm{I}}_{2}$ and the first component of ${\bm{\mu}}_{R}$ and ${\bm{\mu}}_{F}$ being identical. Here, ${\bm{I}}_{2}$ denotes the $2\times 2$ identity matrix. Let $\nabla_{{\bm{x}}}\log p_{\theta}({\bm{x}}|{\bm{c}}_{R})$ be the estimated score function prameterized by $\theta$. We assume that the score estimator is given for each concept as: $\nabla_{\bm{x}}\log p_{\theta}({\bm{x}}|{\bm{c}}_{F})=\nabla_{\bm{x}}\log{\mathcal{N}}({\bm{\mu}}_{F},{\bm{\Sigma}}_{F})=\nabla_{\bm{x}}\log p_{\text{data}}({\bm{x}}|{\bm{c}}_{F})$; and $\nabla_{{\bm{x}}}\log p_{\theta}({\bm{x}}|{\bm{c}}_{R})=\nabla_{{\bm{x}}}\log{\mathcal{N}}(\hat{{\bm{\mu}}}_{R},\hat{{\bm{\Sigma}}}_{R})$, where $\hat{{\bm{\mu}}}_{R}={{\bm{\mu}}}_{R}$ and $\hat{{\bm{\Sigma}}}_{R}=diag(\sigma,1)$. Here, $diag(\sigma,1)$ denotes the $2\times 2$ diagonal matrix with $\sigma$ and 1 as its diagonal elements.
@@ 3.2 Proposed Framework: R2F
[…]
@@ 3.3 Extention with Region-guided Diffusion Models: R2F+
[…]
@@ 4.2 Main Results of R2F
[…]
@@ 4.3 Flexibility across Various Diffusion Models and LLMs
[…]
@@ Appendix C LLM Instruction and Algorithm Pseudocode for R2F+
0:  $\mathbf{c}_{\mathrm{input}}$: an input prompt, $T_{\mathrm{CA}}$: the number of steps for performing cross-attention control, $N_{\mathrm{CA}}$: the number of iterations for performing gradient descent in each cross-attention control step, $\eta_{\mathrm{CA}}$: the coefficient for cross-attention control, $T_{\mathrm{LF}}$: the number of steps for performing latent fusion
0:  $\mathbf{x}_{0}$: an overall image
@@
2:   $\left(\mathbf{c}_{\mathrm{base}},\{(\mathbf{c}^{i},\mathbf{c}^{i}_{R},\mathbf{c}^{i}_{F},V^{i},R^{i})\}_{i=1}^{m}\right)\leftarrow\textsc{GetResponseFromLLM}(\mathbf{c}_{\mathrm{input}})$
@@
[…]
@@
[…]
@@ Appendix D Details for Constructing RareBench
Shape banana-shaped, star-shaped, ax-shaped, butterfly-shaped, oval-shaped, donut-shaped, hand-shaped, gear-shaped, heart-shaped, diamond-shaped
Texture flower-patterned, zebra-striped, tiger-striped, black-white-checkered, made of marble, made of diamonds, made of plastic, made of glass, made of steel, made of cloud
Action dancing, walking, running, crawling, flying, swimming, driving a car, yawning, smiling, crying, cheerleading
@@
[…]
@@ Appendix I Failure Results of T2I-CompBench’s Auto-evaluation Metrics
GPU Time. The time taken for R2F+ is approximately 72 sec, which can be decomposed as 1) masked latent generation via object-wise R2F, which takes around 42 sec, and 2) region-controlled concept guidance, which takes around 30 sec. Specifically, for the process of 1), the generation of each object takes around 20 sec (same as R2F) with an additional 1 sec for masking, resulting in a total of 42 sec for two objects. For the process of 2), the majority of the increased computation time compared to R2F is attributed to attention control, which adds around 10 sec. Consequently, for a prompt with $N$ objects, the time complexity of R2F+ is expected as $N*(T+1)+(T+10)$, where $T$ is the inference time of R2F or SD3.

### v4

@@ Abstract
State-of-the-art text-to-image (T2I) diffusion models often struggle to generate rare compositions of concepts, e.g., objects with unusual attributes. In this paper, we show that the compositional generation power of diffusion models on such rare concepts can be significantly enhanced by the Large Language Model (LLM) guidance. We start with empirical and theoretical analysis, demonstrating that exposing frequent concepts relevant to the target rare concepts during the diffusion sampling process yields more accurate concept composition. Based on this, we propose a training-free approach, R2F, that plans and executes the overall rare-to-frequent concept guidance throughout the diffusion inference by leveraging the abundant semantic knowledge in LLMs. Our framework is flexible across any pre-trained diffusion models and LLMs, and seamlessly integrated with the region-guided diffusion approaches. Extensive experiments on three datasets, including our newly proposed benchmark, RareBench, containing various prompts with rare compositions of concepts, R2F significantly surpasses existing models including SD3.0 and FLUX by up to $28.1\%p$ in T2I alignment. Code is available at link.
Figure 1: Generated images from prompts with rare compositions of concepts (==attribute++object; highlighted in red). These objects possess attributes not typically associated with them, making such combinations difficult to observe. While state-of-the-art pre-trained and LLM-grounded text-to-image diffusion models, SD3.0 (Esser et al., 2024), FLUX (BlackForestLabs, 2024), and RPG (Yang et al., 2024), struggle to generate such concepts, our training-free approach, R2F, exhibits superior results.
@@ 1 Introduction
Recent advancements in text-to-image (T2I) diffusion models have achieved unprecedented success in generating highly realistic and diverse images (Zhang et al., 2023a, Saharia et al., 2022). However, these models often struggle to accurately generate images from rare and complex prompts (Samuel et al., 2024, Parashar et al., 2024). Accordingly, many studies have focused on enhancing their T2I alignment performance (Rassin et al., 2024). Recently, several approaches have further tried to ground large language models (LLMs) into diffusion models, so-called LLM-grounded diffusion models, showing state-of-the-art results by effectively leveraging LLMs’ knowledge in diffusion inference.
@@ 2.3 LLM-grounded Diffusion
LLMs have shown promising abilities in language comprehension (Achiam et al., 2023, Zhao et al., 2023, Wang et al., 2024a). Powered by this advance, recent works have attempted to ground LLMs into diffusion models to provide prior knowledge and conditions for compositional image generation. LayoutGPT (Feng et al., 2024), LMD (Lian et al., 2023), and CompAgent (Wang et al., 2024d) use LLMs to decompose a given prompt into sub-prompts per object and extract their corresponding bounding boxes. RPG (Yang et al., 2024) further adopts recaptioning and planning for complementary regional diffusion. ELLA (Hu et al., 2024) use LLMs to dynamically extract timestep-dependent conditions from intricate prompts. Some work utilizes LLMs in iterative image editing (Wu et al., 2024, Gani et al., 2023, Wang et al., 2024c). While these LLM-grounded diffusion approaches have succeeded in spatial compositions, they still struggle in non-spatial compositions for rare concepts.
@@ 3.1 Compositional Text-to-image Generation for Rare Concepts
Theoretical Motivation. To analyze how relevant frequent concepts help the rare concept composition, we consider a simple setting where two texts are given: one with a rare concept ${\bm{c}}_{R}$ (e.g.,“furry frog”), and another with a frequent one ${\bm{c}}_{F}$ (e.g., “furry dog”). Let ${\bm{x}}\in\mathbb{R}^{2}$ be an image representation, where the first and second dimension represents the attribute (e.g., “furry”) and object (e.g., “animal”), respectively. We assume the ground truth conditional distributions $p_{\text{data}}({\bm{x}}|{\bm{c}}_{R})$ and $p_{\text{data}}({\bm{x}}|{\bm{c}}_{F})$ follow Gaussian distributions ${\mathcal{N}}({\bm{\mu}}_{R},{\bm{\Sigma}}_{R})$ and ${\mathcal{N}}({\bm{\mu}}_{F},{\bm{\Sigma}}_{F})$ (Liang et al., 2024), where ${\bm{\Sigma}}_{R}={\bm{\Sigma}}_{F}={\bm{I}}_{2}$ and the first component of ${\bm{\mu}}_{R}$ and ${\bm{\mu}}_{F}$ being identical. Here, ${\bm{I}}_{2}$ denotes the $2\times 2$ identity matrix. Let $\nabla_{{\bm{x}}}\log p_{\theta}({\bm{x}}|{\bm{c}}_{R})$ be the estimated score function prameterized by $\theta$. We assume that the score estimator is given for each concept as: $\nabla_{\bm{x}}\log p_{\theta}({\bm{x}}|{\bm{c}}_{F})=\nabla_{\bm{x}}\log{\mathcal{N}}({\bm{\mu}}_{F},{\bm{\Sigma}}_{F})=\nabla_{\bm{x}}\log p_{\text{data}}({\bm{x}}|{\bm{c}}_{F})$; and $\nabla_{{\bm{x}}}\log p_{\theta}({\bm{x}}|{\bm{c}}_{R})=\nabla_{{\bm{x}}}\log{\mathcal{N}}(\hat{{\bm{\mu}}}_{R},\hat{{\bm{\Sigma}}}_{R})$, where $\hat{{\bm{\mu}}}_{R}={{\bm{\mu}}}_{R}$ and $\hat{{\bm{\Sigma}}}_{R}=diag(\sigma,1)$. Here, $diag(\sigma,1)$ denotes the $2\times 2$ diagonal matrix with $\sigma$ and 1 as its diagonal elements.
@@ 3.2 Proposed Framework: R2F
Rare-to-Frequent Concept Mapping. To extract proper concept mapping, we use LLM instruction with chain-of-thought prompting (See Appendix B for the full LLM instruction and in-context learning examples). Given a text prompt $\mathbf{c}$}, R2F performs the concept mapping process step by step: (1) It decomposes the prompt into $m$ sub-prompts per object, ${\mathbf{c}}=\{{\mathbf{c}}^{i}\}^{m}_{i=1}$. For each sub-prompt ${\mathbf{c}}^{i}$, (2) it identifies rare concepts and (3) extracts a frequent concept contextually relevant to the rare concept, making a rare-to-frequent concept mapping $\hat{{\mathbf{c}}}^{i}=({\mathbf{c}}^{i},{\mathbf{c}}^{i}_{R},{\mathbf{c}}^{i}_{F})$. Specifically, to find the frequent concept ${\mathbf{c}}^{i}_{F}$, it splits each rare concept into the main noun object and its attributes and then replaces the main noun object into another object contextually more fit to the attributes. Additionally, (4) It extracts a visual detail level $V^{i}$ (an integer value from 1 to 5) needed to draw each rare concept for determining an appropriate stop point of concept guidance, based on prior observation (Hertz et al., 2022) that rough visual features (e.g., shape) are highly affected by diffusion latents at early sampling steps while detailed visual features (e.g., texture) are influenced at the later sampling steps. The final mapped rare-to-frequent concept output is formed as $\hat{{\mathbf{c}}}=\{({\mathbf{c}}^{i},{\mathbf{c}}^{i}_{R},{\mathbf{c}}^{i}_{F},V^{i})\}_{i=1}^{m}$.
@@ 3.3 Extention with Region-guided Diffusion Models: R2F+
Region-controlled Alternating Concept Guidance. With the obtained per-object rare-to-frequent prompts and their masked latents, we finally generate the whole objects all at once via region-controlled alternating concept guidance. To force each object to appear in the proper region, we adopt two popular region control techniques: (1) cross-attention control (Chen et al., 2024c), and (2) latent fusion (Lian et al., 2023). At each step $t$, the latent $\mathbf{z}_{t}$ is first modified by the cross-attention control: ${\mathbf{z}}_{t}^{\prime}\xleftarrow{}{\mathbf{z}}_{t}-\eta\nabla_{{\mathbf{z}}_{t}}\sum_{i=1}^{m}E(A^{i},R^{i})\quad\textit{(cross-attention control)},$ where $A^{i}$ is an average cross-attention map for the object $i$ across the diffusion layers, $E(A,R)=\big(1-\sum_{(x,y)\in R}A_{x,y}/\sum_{(x,y)}A_{x,y}\big)^{2}$ is an energy function, and $\eta$ is a controlling parameter. Next, the original diffusion sampling is performed, as ${\mathbf{z}}_{t-1}\leftarrow p_{\theta}({{\mathbf{z}}}_{t}^{\prime},\tilde{\mathbf{c}}^{t})$. Afterward, we compound the controlled latent ${\mathbf{z}}_{t-1}$ with every object-wise latent $\bar{\mathbf{z}}^{i}_{t-1}$ sequentially via the latent fusion: ${{\mathbf{z}}}_{t-1}\leftarrow{\mathbf{z}}_{t-1}\odot(1-M^{i})+\bar{{\mathbf{z}}}^{i}_{t-1}\odot M^{i}\quad\textit{(latent fusion)}.$ Both the cross-attention control and the latent fusion are applied only to initial sampling steps until the stop points, integrated with the alternating rare-to-frequent concept guidance.
@@ 4.2 Main Results of R2F
Models Single Object Multi Objects Property Shape Texture Action Complex Concat Relation Complex SDXL 60.0 56.9 71.3 47.5 58.1 39.4 35.0 47.5 LMD 23.8 35.6 27.5 23.8 35.6 33.1 34.4 33.1 RPG 33.8 54.4 66.3 31.9 37.5 21.9 15.6 9.4 ELLA 31.3 63.6 64.4 43.1 66.3 42.5 50.6 51.9 $\text{{R2F}{}}_{\text{sdxl}}$ 71.3 71.9 73.8 54.4 70.6 50.6 36.0 52.8 IterComp 63.8 66.9 61.3 65.6 61.9 41.3 29.4 53.1 $\text{{R2F}{}}_{\text{itercomp}}$ 78.1 77.5 79.4 66.9 63.9 41.5 36.6 53.4 SD3.0 49.4 76.3 53.1 71.9 65.0 55.0 51.2 70.0 $\text{{R2F}{}}_{\text{sd3.0}}$ 89.4 79.4 81.9 80.0 72.5 70.0 58.8 73.8
Figure 7: Generated images by $\text{{R2F}{}}_{\text{sdxl}}$.
@@ 4.3 Flexibility across Various Diffusion Models and LLMs
Effectiveness across Different Diffusion Models. Table 4.2 shows GPT-4o evaluated T2I alignment performance of R2F combined with various diffusion backbones, including SDXL, IterComp, and SD3.0. R2F consistently improves the performance of three backbones in all cases, regardless of the backbone used. Also, $\text{{R2F}{}}_{\text{sdxl}}$ outperforms existing LLM-grounded diffusion baselines based on SDXL in most cases, indicating its effectiveness in leveraging relevant frequent concepts in generating images of rare concepts. Figure 7 visualizes the generated images by $\text{{R2F}{}}_{\text{sdxl}}$.
@@ Appendix C LLM Instruction and Algorithm Pseudocode for R2F+
[…]
@@
[…]
@@
26:function $\textsc{GetScheduledPrompt}\left(t,\mathbf{c},\{(\mathbf{c}^{i}_{R},\mathbf{c}^{i}_{F},V^{i})\}_{i=1}^{k}\right)$
@@
32:function $\textsc{CrossAttentionControl}\left(\mathbf{z}_{t},\{(\mathbf{c}^{i},R^{i})\}_{i=1}^{k}\right)$
@@ Appendix D Details for Constructing RareBench
Shape
banana-shaped, star-shaped, ax-shaped, butterfly-shaped, oval-shaped, donut-shaped,
hand-shaped, gear-shaped, heart-shaped, diamond-shaped
Texture
flower-patterned, zebra-striped, tiger-striped, black-white-checkered, made of marble,
made of diamonds, made of plastic, made of glass, made of steel, made of cloud
Action
dancing, walking, running, crawling, flying, swimming, driving a car, yawning, smiling,
crying, cheerleading
@@
Appendix E GPT Instruction for Calculating $\%$ Rareness
Since real-world T2I datasets typically exhibit a long-tailed nature (Xu et al., 2023, Park et al., 2020), they contain many rare concepts (in the tail distributions). To measure whether a prompt contains rare concepts that are very difficult to observe in the real-world, we ask GPT4 with the yes or no binary question using the following instructions. “You are an assistant to evaluate if the text prompt contains rare concepts that exist infrequently or not in the real world. Evaluate if rare concepts are contained in the text prompt: PROMPT, The answer format should be YES or NO, without any reasoning.”. Formally, the $\%$ rareness of each test dataset ${\mathcal{C}}_{\text{test}}$ is calculated as $\%\text{Rareness}({\mathcal{C}}_{\text{test}})=1/|{\mathcal{C}}_{\text{test}}|\sum_{{\bm{c}}\in{\mathcal{C}}_{\text{test}}}{\mathbbm{1}(\text{GPT}_{\text{rare}}({\bm{c}})==\text{Yes})}$, where $\text{GPT}_{\text{rare}}({\bm{c}})$ is the binary answer of rareness from GPT.
@@ Appendix I Failure Results of T2I-CompBench’s Auto-evaluation Metrics
GPU Time. The time taken for R2F+ is approximately 72 sec, which can be decomposed as 1) masked latent generation via object-wise R2F, which takes around 42 sec, and 2) region-controlled concept guidance, which takes around 30 sec. Specifically, for the process of 1), the generation of each object takes around 20 sec (same as R2F) with an additional 1 sec for masking, resulting in a total of 42 sec for two objects. For the process of 2), the majority of the increased computation time compared to R2F is attributed to attention control, which adds around 10 sec. Consequently, for a prompt with $N$ objects, the time complexity of R2F+ is expected as N(T+1)+(T+10)N*(T+1)+(T+10), where $T$ is the inference time of R2F or SD3.